Can you factor this so I can check my work and did it properly please? I really appreciate your help.:
x^2 - 9x = -8

Answer by Nate(3500)

x^2 - 9x + 8 = 0
x^2 - 8x - x + 8 = 0
(x^2 - 8x) + (-x + 8) = 0
x(x - 8) - (x - 8) = 0
(x - 1)(x - 8) = 0
x = 1, x = 8